# MIVE: New Design and Benchmark for Multi-Instance Video Editing

### v1

@@ Abstract
Recent AI-based video editing has enabled users to edit videos through simple text prompts, significantly simplifying the editing process. However, recent zero-shot video editing techniques primarily focus on global or single-object edits, which can lead to unintended changes in other parts of the video. When multiple objects require localized edits, existing methods face challenges, such as unfaithful editing, editing leakage, and lack of suitable evaluation datasets and metrics. To overcome these limitations, we propose a zero-shot Multi-Instance Video Editing framework, called MIVE. MIVE is a general-purpose mask-based framework, not dedicated to specific objects (e.g., people). MIVE introduces two key modules: (i) Disentangled Multi-instance Sampling (DMS) to prevent editing leakage and (ii) Instance-centric Probability Redistribution (IPR) to ensure precise localization and faithful editing. Additionally, we present our new MIVE Dataset featuring diverse video scenarios and introduce the Cross-Instance Accuracy (CIA) Score to evaluate editing leakage in multi-instance video editing tasks. Our extensive qualitative, quantitative, and user study evaluations demonstrate that MIVE significantly outperforms recent state-of-the-art methods in terms of editing faithfulness, accuracy, and leakage prevention, setting a new benchmark for multi-instance video editing. The project page is available at https://kaist-viclab.github.io/mive-site/.
Figure 1: Given a video, instance masks, and target instance captions, our MIVE framework enables faithful and disentangled edits in (a) single- and (b)-(c) multi-instance levels, as well as an applicability to more fine-grained (d) partial instance level without the need for additional training. Unlike previous methods, our MIVE does not rely on global edit captions, but leverages individual instance captions. Each object mask is color-coded to match its corresponding edit caption. Zoom-in for better visualization.
@@ 1 Introduction
Recently, Ground-A-Video (GAV) [25] demonstrated that simultaneous multi-object VE is feasible using a T2I model [44] fine-tuned on grounding conditions [29]. However, GAV still suffers from attention leakage [61], especially when the objects’ bounding boxes overlap (e.g., the statue also gaining flower crown in Fig. 2-(b)). EVA [63], a concurrent work, attempts to address the leakage by assigning edit prompts to object masks. However, it is designed for human-centric videos [63] and does not demonstrate diverse object (e.g., vehicles) editing capability.
@@
Furthermore, both GAV [25] and EVA [63] have been tested on limited datasets, insufficient for comprehensive testing of multi-instance VE across diverse viewpoints, instance sizes, and numbers of instances. The global editing metrics used in these methods further fail to accurately measure local editing quality, essential for multi-instance VE.
@@
From the above, four critical challenges persist in multi-instance VE: (i) Attention leakage. The absence of local control in pre-trained T2I models, imprecise input conditions [25], and the use of a single global edit caption [70] hinder previous methods from effectively disentangling edit prompts; (ii) Lack of versatile multi-instance VE methods. While EVA [63] exhibits per-object editing capabilities, it is not a general-purpose VE framework; (iii) Unfaithful editing. The lack of techniques to enhance editing faithfulness in multi-instance VE tasks often results in inaccurate editing; (iv) Lack of evaluation dataset and metrics. Recent methods [63, 25] are tested on limited datasets using metrics inadequate for evaluating local VE quality.
@@
To overcome these, we present MIVE, a general-purpose zero-shot Multi-Instance Video Editing framework to achieve faithful edits and reduce attention leakage by disentangling multi-instance edits. MIVE integrates easily into existing T2I models and achieves multi-instance editing capabilities through two key modules: (i) Inspired by prior works [55, 46], we design Disentangled Multi-instance Sampling (DMS) which significantly reduces attention leakage; and (ii) To enhance editing faithfulness and increase the likelihood of objects appearing within their masks, we introduce Instance-centric Probability Redistribution (IPR) in the cross-attention layers. Moreover, we evaluate MIVE on our proposed dataset, called MIVE Dataset, of 200 videos using standard metrics and a novel metric, the Cross-Instance Accuracy (CIA) Score, to quantify attention leakage. Our contributions are as follows:
@@
  • We propose a novel general-purpose mask-based zero-shot Multi-Instance Video Editing framework, called MIVE, which enables multi-instance editing for videos;
@@
  • We propose to disentangle multi-instance video editing through our (i) Disentangled Multi-instance Sampling (DMS) that reduces editing leakage and (ii) Instance-centric Probability Redistribution (IPR) that enhances editing localization and faithfulness;
@@
  • Our extensive experiments validate the effectiveness of our MIVE in disentangling multiple edits across various instances and achieving faithful editing, significantly outperforming the most recent SOTA video editing methods.
Figure 3: The overall framework of Multi-Instance Video Editing (MIVE). Our Disentangled Multi-instance Sampling (DMS, Sec. 3.2) consists of latent parallel sampling (LPS, blue box), latent fusion (yellow box) to fuse different instance latents, re-inversion (red box) to harmonize the latents after fusion, and noise parallel sampling (NPS, green box). In addition, our Instance-centric Probability Redistribution (IPR, Sec. 3.3) provides better spatial control.
@@ 2 Related Works
Zero-shot text-guided video editing. Recent advancements in diffusion models [20, 49, 50] have accelerated the evolution of both text-to-image (T2I) [42, 38, 45, 12, 44] and text-to-video (T2V) [4, 67, 57, 21, 17, 15, 53] models for generative tasks. These breakthroughs have led to the development of numerous video editing (VE) frameworks, where pre-trained models serve as backbones. Most VE methods [40, 58, 16, 11, 26, 25, 70] rely on pre-trained T2I models [44, 29, 69], as T2V models remain less accessible to the public [5] or are computationally expensive to run [53]. To facilitate VE, recent methods have either fine-tuned T2I models on large video datasets [71, 56, 15], optimized on a single input video [58, 32], or leveraged zero-shot inference [70, 11, 26, 62, 16, 25, 40]. Our work falls within zero-shot VE, allowing editing without additional training.
Zero-shot VE methods often prioritize temporal consistency through various techniques such as modifying the U-Net’s self-attention layers [58, 54, 11, 62, 6], applying optimization strategies [62, 6], smoothing input latents [25, 70], linearly combining features [16], or transforming input latents [26]. Our approach focuses more on achieving faithful editing and reducing attention leakage while adopting a previous method [11] to achieve temporal consistency.
@@
Local video editing and techniques from image generation. Similar to image generation and editing, the need for fine-grained control in videos has led to the development of several local VE methods [71, 32, 25, 63]. These methods usually adopt spatial control techniques from image generation such as training additional adapters [69, 1, 29, 64, 55] or employing zero-shot techniques, e.g., optimization [3, 8, 9, 60], attention modulation [28], and multi-branch sampling [46]. In particular, AVID [71] uses masks and retrains the Stable Diffusion (SD) inpainting model [44], while Video-P2P [32] leverages an attention control method [18], enabling both to achieve finer control.
@@
Extending the local control for multi-instance VE scenarios is relatively underexplored, with few works [25, 63] tackling this challenge by similarly adopting image generation techniques [29, 28] to localize multi-instance VE. GAV [25] leverages GLIGEN [29], a framework that uses bounding boxes and integrates gated self-attention layers within the transformer blocks, allowing spatial control. However, GLIGEN requires retraining the gated self-attention layers for every new T2I model (e.g. SDv1.5 and SDv2.1), reducing its flexibility. EVA [63] uses masks and discrete text embeddings for spatial control, alongside attention modulation [28] for the spatio-temporal attention (STA). However, inaccurate feature correspondences [65] in the STA amplify flickering since attention modulation, if not controlled well, causes artifacts [28]. Our work similarly modulates attention values but only within the cross-attention layers, enabling spatial control without affecting the STA.
@@ 3.1 Overall Framework
Figure 3 illustrates the overall framework of MIVE. MIVE falls under the category of inversion-based VE [40] (see Supplemental Sec. A for preliminaries). Starting with the initial latents $z_{0}^{1:N}=\mathcal{E}(f^{1:N})$ generated using the VAE encoder $\mathcal{E}$ [44], we employ the non-inflated U-Net to invert the latents using DDIM Inversion [49]. This yields a sequence of inverted latents, $\{\boldsymbol{\tilde{z}_{t}}\}_{t=0}^{T}=\{\tilde{z}_{t}^{1:N}\}_{t=0}^{T}$, that we store for later use with $T$ representing the number of denoising steps.
For the sampling, we introduce the Disentangled Multi-instance Sampling (DMS, Sec. 3.2), inspired by image generation methods [55, 46], to disentangle the multi-instance VE process and minimize attention leakage. Each instance is independently modified using latent parallel sampling (blue box), and the multiple denoised instance latents are harmonized through noise parallel sampling preceded by latent fusion and re-inversion (green, yellow, and red boxes, respectively). We employ the 3D U-Net [58] and 3D ControlNet [69] following [25, 70] during sampling, with depth maps $\boldsymbol{d}=d^{1:N}_{i}$ obtained via MiDas [43] as conditions.
@@
Our DMS requires each edited instance $i$ to appear within its mask $\boldsymbol{m_{i}}$. Vanilla cross-attention [44] can not ensure the target edit stays within $\boldsymbol{m_{i}}$ (see Fig. 7-(b)). To address this, we introduce Instance-centric Probability Redistribution (IPR, Sec. 3.3) in the cross-attention, boosting the accuracy of edit placement inside $\boldsymbol{m_{i}}$. Finally, we employ FLATTEN [11] to ensure temporal consistency.
3.2 Disentangled Multi-instance Sampling
To disentangle the multi-instance VE process and reduce attention leakage, we propose our Disentangled Multi-instance Sampling (DMS). As shown in Fig. 3, DMS comprises of two sampling strategies: (1) Latent Parallel Sampling (LPS) shown in blue box and (2) Noise Parallel Sampling (NPS) shown in green box preceded by latent fusion (yellow box) and re-inversion (red box). In the LPS, we independently edit each instance $i$ using its target caption $c_{i}$ and its mask $\boldsymbol{m_{i}}$ through DDIM [49] denoising as follows:
[…]
Then, we fuse the denoised instance latents $\{\hat{z}_{t-1}^{1:N}\}_{i}$ with the inverted latents $\tilde{z}_{t-1}^{1:N}$ to obtain the edited instance latents $\{\bar{z}_{t-1}^{1:N}\}_{i}$ using masking as:
$\{\bar{z}_{t-1}^{1:N}\}_{i}=\{\hat{z}_{t-1}^{1:N}\}_{i}\cdot\boldsymbol{m_{i}}+\tilde{z}_{t-1}^{1:N}\cdot(1-\boldsymbol{m_{i}}).$ (2)
LPS requires that the edited instance $i$ appears within $\boldsymbol{m_{i}}$ since we are replacing the background with the inverted latents at each step. This is achieved through our proposed IPR (Sec. 3.3), providing the necessary spatial control.
While we can perform multi-instance VE by solely using LPS until the end of denoising, the resulting instances have artifacts because they are sampled independently (see Fig. 6-(c)). Thus, we propose to fuse the multiple instance latents $\{\dot{z}_{t}^{1:N}\}$ obtained from LPS during intermediate sampling steps as follows:
$\dot{z}_{t}^{1:N}=\Sigma_{i=1}^{M}(\{\bar{z}_{t}^{1:N}\}_{i}\cdot\boldsymbol{m_{i}})+\tilde{z}_{t}^{1:N}\cdot\boldsymbol{m_{B}}.$ (3)
where $\boldsymbol{m_{B}}=\boldsymbol{1}-\Sigma_{i=1}^{M}\boldsymbol{m_{i}}$ denotes a background mask. We then perform NPS (details in next paragraph) to yield the intermediate latents $\hat{z}_{t-1}^{1:N}$. We can use $\hat{z}_{t-1}^{1:N}$ and continue alternating between LPS and latent fusion with NPS until the end of denoising. However, the results still contain artifacts, e.g., noise and blur (see Fig. 6-(d)). Thus, we propose performing re-inversion using DDIM for $l$ steps after the latent fusion to obtain $\ddot{z}_{t+l}^{1:N}$ followed by NPS from timestep $t+l$ to $t$ to yield the final fused latent $\ddot{z}_{t}^{1:N}$. Unlike the initial inversion, our re-inversion uses the 3D U-Net and ControlNet to avoid causing additional artifacts (see Supp.).
The goal of our NPS is to harmonize the independent instance latents obtained from the LPS. In this sampling strategy, we use the re-inverted fused latents $\ddot{z}_{t}^{1:N}$ to estimate each instance noise $\hat{n}_{i}^{1:N}=\epsilon_{\theta}(\ddot{z}_{t}^{1:N},c_{i},\boldsymbol{m_{i}},e,t)$ using the instance caption $c_{i}$. We also estimate the noise $\tilde{n}^{1:N}=\epsilon_{\theta}(\tilde{z}_{t}^{1:N},c_{\varnothing},t)$ for the background using the inverted latents $\tilde{z}_{t}^{1:N}$ and an empty caption $c_{\varnothing}$. We then combine $\hat{n}_{i}^{1:N}$ and $\tilde{n}^{1:N}$ to a single noise through masking, and perform one DDIM step to obtain the latents $\hat{z}_{t-1}^{1:N}$:
$\hat{z}_{t-1}^{1:N}=\text{DDIM}(\hat{z}_{t}^{1:N},\Sigma_{i=1}^{M}(\hat{n}_{i}^{1:N}\cdot\boldsymbol{m_{i}})+\tilde{n}^{1:N}\cdot\boldsymbol{m_{B}},t)$ (4)
@@ 3.3 Instance-centric Probability Redistribution
Our sampling requires the edited objects to appear within their masks, highlighting the need for spatial control in the U-Net because the vanilla cross-attention [44] struggles to localize edits (see Fig. 7-(b)). To address this, we propose Instance-centric Probability Redistribution (IPR), inspired by attention modulation [28]. Our IPR further achieves faithful editing with less artifacts, unlike [28] that uses maximization-minimization, potentially disrupting the spatial distribution of initial attention scores and causing artifacts (see Fig. 7-(c)). Fig. 4 shows our IPR (bottom part) and a comparative illustration (top part) versus others.
@@
In our IPR, we split $A_{i,j}$ of an instance into two sets depending on the mask $m$: outside the instance $\boldsymbol{A}_{\boldsymbol{O},j}=\{A_{i,j};m_{i}=0\}$ and inside the instance $\boldsymbol{A}_{\boldsymbol{I},j}=\{A_{i,j};m_{i}=1\}$. Our IPR is based on several experimental observations (see Supp. Sec. E.2 for details): (i) Manipulating the attention probabilities of the padding tokens $A_{i,j\in P}$ may lead to artifacts, so we avoid altering these; (ii) Increasing $S$ token’s probability decreases editing faithfulness, while decreasing it and reallocating those values to the $\boldsymbol{T}$ and $E$ tokens increases editing faithfulness. Thus, for $\boldsymbol{A}_{\boldsymbol{O},j}$, we zero out $\boldsymbol{T}$ and $E$ tokens’ probabilities, reallocating them to $S$ to prevent edits outside the mask (blue dashed box in Fig. 4, bottom). In addtition, we redistribute the attention probability for the object region inside the mask by reducing the value of $S$ and redistributing it to $\boldsymbol{T}$ and $E$ (red dashed box in Fig. 4, bottom). We reduce $\boldsymbol{A}_{\boldsymbol{I},j=S}$ by $\lambda_{S}$, which decays linearly to 0 from $t=T$ to $t=1$, formulated as:
@@
where $W$ is a warm-up value that is linearly decayed from $\lambda$ to 0 in the early sampling steps $t<0.1T$ to increase the editing fidelity especially for small and difficult objects. Increasing $\lambda$ tends to enhance faithfulness but may introduce artifacts. We empirically found that $\lambda=0.4$ is the best trade-off between faithfulness and artifact-less results. Then, we redistribute $\lambda_{S}$ to $\boldsymbol{T}$ and $E$ with the ratios of $\lambda_{r}$ and $1-\lambda_{r}$, respectively. The attention probabilities, $\boldsymbol{A}_{\boldsymbol{I},j\in\boldsymbol{T}}$ and $\boldsymbol{A}_{\boldsymbol{I},j=E}$, of $\boldsymbol{T}$ and $E$ are updated as follows:
@@
where $N_{\boldsymbol{T}}$ denotes the number of text tokens. Increasing $\lambda_{r}$ may lead to enhancing the editing details for certain tokens, while decreasing it may increase the overall editing fidelity. $\lambda_{r}=0.6$ is empirically set for all experiments.

### v2

# PRIMEdit: Probability Redistribution for Instance-aware Multi-object Video Editing with Benchmark Dataset

@@ Abstract
Recent AI-based video editing has enabled users to edit videos through simple text prompts, significantly simplifying the editing process. However, recent zero-shot video editing techniques primarily focus on global or single-object edits, which can lead to unintended changes in other parts of the video. When multiple objects require localized edits, existing methods face challenges, such as unfaithful editing, editing leakage, and lack of suitable evaluation datasets and metrics. To overcome these limitations, we propose Probability Redistribution for Instance-aware Multi-object Video Editing (PRIMEdit). PRIMEdit is a zero-shot framework that introduces two key modules: (i) Instance-centric Probability Redistribution (IPR) to ensure precise localization and faithful editing and (ii) Disentangled Multi-instance Sampling (DMS) to prevent editing leakage. Additionally, we present our new MIVE Dataset for video editing featuring diverse video scenarios, and introduce the Cross-Instance Accuracy (CIA) Score to evaluate editing leakage in multi-instance video editing tasks. Our extensive qualitative, quantitative, and user study evaluations demonstrate that PRIMEdit significantly outperforms recent state-of-the-art methods in terms of editing faithfulness, accuracy, and leakage prevention, setting a new benchmark for multi-instance video editing.
Figure 1: Given a video, instance masks, and target instance captions, our PRIMEdit framework enables faithful and disentangled edits in (a) single- and (b)-(c) multi-instance levels, as well as an applicability to more fine-grained (d) partial instance level without the need for additional training. Unlike previous methods, our PRIMEdit does not rely on global edit captions, but leverages individual instance captions. Each object mask is color-coded to match its corresponding edit caption. Zoom-in for better visualization.
@@ 1 Introduction
Recently, Ground-A-Video (GAV) [28] demonstrated that simultaneous multi-object VE is feasible using a T2I model [46] fine-tuned on grounding conditions [33]. However, GAV suffers from attention leakage [62], especially when the objects’ bounding boxes overlap (e.g., the statue also gaining flower crown in Fig. 2-(b)). VideoGrain [65] attempts to address the leakage by instead using object masks and attention modulation [32]. However, VideoGrain’s use of attention modulation in the spatio-temporal attention (STA) imposes significant memory and computation bottlenecks, particularly as the number of video frames increases see Supplemental). VideoGrain also relies on global edit captions to describe the edited scenes, which may be unavailable or challenging to generate in scenes with a large number of objects.
@@
Furthermore, both GAV [28] and VideoGrain [65] have been tested on limited datasets, insufficient for comprehensive testing of multi-instance VE across diverse viewpoints, instance sizes, and numbers of instances. The global editing metrics used in these methods further fail to precisely measure local editing quality, essential for multi-instance VE.
@@
From the above, four critical challenges persist in multi-instance VE: (i) Attention leakage. The absence of local control in pre-trained T2I models, the imprecise input conditions [28] and the use of a single global edit caption [72] hinder previous methods from effectively disentangling edit prompts; (ii) Unfaithful editing. The lack of techniques to enhance editing faithfulness in multi-instance VE tasks often results in inaccurate editing; (iii) Lack of fast and intuitive multi-instance VE methods. While modulation in [65] enables per-object editing, it slows down the generative editing process. Moreover, its reliance on full-scene descriptions makes editing cumbersome, especially in highly cluttered videos. (iv) Lack of evaluation dataset and metrics. Recent methods [65, 28] are tested on limited datasets using metrics inadequate for evaluating local VE quality.
@@
To overcome these, we present our Probability Redistribution for Instance-aware Multi-object Video Editing framework (PRIMEdit), a more intuitive zero-shot multi-instance VE method to achieve faithful edits and reduce attention leakage by disentangling multi-instance edits. Our PRIMEdit can easily adopt existing T2I and T2V models into a unified framework and achieve multi-instance editing capabilities through two key modules: (i) To enhance editing faithfulness and increase the likelihood of objects appearing within their masks, we introduce Instance-centric Probability Redistribution (IPR, Sec. 3.2) in the cross-attention layers; and (ii) Inspired by prior works [55, 48], we design Disentangled Multi-instance Sampling (DMS, Sec. 3.3) which can significantly reduce attention leakage. Moreover, we evaluate PRIMEdit on our newly proposed benchmark dataset, called MIVE Dataset, of 200 videos using standard metrics and a novel metric, the Cross-Instance Accuracy (CIA) Score, to quantify attention leakage. Our contributions are as follows:
@@
  • We propose a novel zero-shot multi-instance video editing framework, called PRIMEdit, which enables fast and intuitive multi-instance editing for videos;
@@
  • We propose to disentangle multi-instance video editing through our (i) Instance-centric Probability Redistribution (IPR) that enhances editing localization and faithfulness and (ii) Disentangled Multi-instance Sampling (DMS) that reduces editing leakage;
@@
  • Our extensive experiments validate the effectiveness of our PRIMEdit in disentangling multiple edits across various instances and achieving faithful editing with better temporal consistency, significantly outperforming the most recent SOTA video editing methods.
Figure 3: The overall framework of our PRIMEdit, given $M$ number of multi-instance captions $c_{i}$ with corresponding instance masks $\boldsymbol{m_{i}}$ for editing. Our Disentangled Multi-instance Sampling (DMS, Sec. 3.3) consists of series noise sampling (SNS, yellow box), latent fusion (green box) to fuse different instance latents, re-inversion (purple box) to harmonize the latents after fusion, and parallel noise sampling (PNS, blue box). In addition, our Instance-centric Probability Redistribution (IPR, Sec. 3.2) provides better spatial control.
@@ 2 Related Works
Zero-shot text-guided video editing. Recent advancements in diffusion models [25, 51, 52] have accelerated the evolution of both text-to-image (T2I) [44, 40, 47, 15, 46] and text-to-video (T2V) [4, 69, 57, 26, 22, 18, 54] models for generative tasks. These breakthroughs led to the development of numerous video editing (VE) frameworks, where pre-trained models serve as backbones. Most VE methods [42, 59, 20, 14, 30, 28, 72] rely on pre-trained T2I models [46, 33, 71], as early T2V models were either not publicly accessible [5] or computationally expensive to run [54]. To facilitate VE, recent methods have either fine-tuned T2I models on large video datasets [73, 56, 18], optimized on a single input video [59, 36], or leveraged zero-shot inference [72, 14, 30, 63, 20, 28, 42, 65]. Our work falls within zero-shot VE, allowing editing without additional training.
@@
Local video editing based on image generation and editing. Similar to image generation and editing, the need for fine-grained control in videos has led to the development of several local VE methods [73, 36, 28, 65]. These methods usually adopt spatial control techniques from image generation such as training additional adapters [71, 1, 33, 66, 55] or employing zero-shot techniques, e.g., attention manipulation [32, 23], optimization [3, 9, 10, 61], and multi-branch sampling [48]. In particular, AVID [73] uses masks and retrains the Stable Diffusion (SD) inpainting model [46], while Video-P2P [36] leverages an attention control method [23], enabling both to achieve finer control.
@@
Extending the local control for multi-instance VE scenarios is relatively underexplored, with few works [28, 65] tackling this challenge by similarly adopting image generation techniques [33, 32] to localize VE. GAV [28] leverages GLIGEN [33] and integrates gated self-attention layers within the transformer blocks, allowing for spatial control. However, GLIGEN requires retraining the gated self-attention layers for every new T2I model (e.g. SDv1.5 and SDv2.1), reducing its flexibility. VideoGrain [65] uses masks and attention modulation [32] for the spatio-temporal attention (STA) and cross-attention layers. However, manipulating the STA increases computational costs as the number of edited frames increases. Our work introduces a novel redistribution technique that manipulates values only within the cross-attention layers, allowing for seamless integration with any T2I or T2V model and spatial control with minimal computational overhead.
Text-to-video models. A growing trend in T2V generation is training a Diffusion Transformer (DiT) [5, 11] on largescale video datasets. Prior to DiTs, T2V models were built by adding temporal layers to pre-trained T2I models [4, 19, 21] and fine-tuning them on video datasets for temporal modeling as it is more efficient.
Recently, several works [73, 67] have started adopting fine-tuned T2V models in VE to achieve better temporal consistency. AVID [73] extends a pre-trained T2I model [46] to T2V by inserting motion modules and finetunes on a large video dataset to enable fine-grained video editing. DMT [67] introduced a space-time feature loss derived from a T2V model [54, 6] to transfer motion from objects in a reference video to target edit objects in a zero-shot manner. Similarly, we use a pre-trained publicly available T2V model [21] to achieve better temporal consistency.
@@ 3.1 Overall Framework
Fig. 3 illustrates the overall framework of our PRIMEdit. PRIMEdit falls under the category of inversion-based VE [42] (see Supp. for preliminaries). Starting with the initial latents $z_{0}^{1:N}=\mathcal{E}(f^{1:N})$ generated using the VAE encoder $\mathcal{E}$ [46], we employ the 3D U-Net [46, 21] to invert the latents using DDIM Inversion [51]. This yields a sequence of inverted latents, $\{\boldsymbol{\tilde{z}_{t}}\}_{t=0}^{T}=\{\tilde{z}_{t}^{1:N}\}_{t=0}^{T}$, that we store for later use with $T$ representing the number of denoising steps.
@@
The vanilla cross-attention within the U-Net [46] cannot ensure the target edit to stay within $\boldsymbol{m_{i}}$ (see Fig. 6-(b)). To address this, we introduce Instance-centric Probability Redistribution (IPR, Sec. 3.2) in the cross-attention, boosting the accuracy of edit placement inside $\boldsymbol{m_{i}}$.
For the sampling, we introduce the Disentangled Multi-instance Sampling (DMS, Sec. 3.3), inspired by image generation methods [55, 48], to disentangle the multi-instance VE process, and minimize attention leakage. Each instance is independently modified using Series Noise Sampling (SNS, blue box in Fig. 3), and the multiple denoised instance latents are harmonized through Parallel Noise Sampling (PNS) preceded by Latent Fusion and Re-inversion (green, yellow, and red boxes, respectively, in Fig. 3). We employ ControlNet [71] during sampling, with depth maps $\boldsymbol{d}=d^{1:N}_{i}$ obtained via MiDas [45] as conditions.
@@ 3.2 Instance-centric Probability Redistribution
Our sampling requires the edited objects to appear within their masks, highlighting the need for spatial control in the U-Net because the vanilla cross-attention [46] struggles to localize edits (see Fig. 6-(b)). To address this, we propose Instance-centric Probability Redistribution (IPR), inspired by attention modulation [32]. Our IPR enables faithful editing while operating exclusively within the cross-attention layers, avoiding the significant computational overhead associated with using [32] in both the spatio-temporal and cross-attention layers [65] to produce high-quality results (see Fig. 6-(c)). Fig. 4 shows our IPR (bottom part) and a comparative illustration (top part) versus others.
@@
In our IPR, we split $A_{i,j}$ of an instance into two sets depending on the mask $m$: outside the instance $\boldsymbol{A}_{\boldsymbol{O},j}=\{A_{i,j};m_{i}=0\}$ and inside the instance $\boldsymbol{A}_{\boldsymbol{I},j}=\{A_{i,j};m_{i}=1\}$. Our IPR is based on several experimental observations (see Supp. for details): (i) Manipulating the attention probabilities of the padding tokens $A_{i,j\in P}$ may lead to artifacts, so we avoid altering these; (ii) Increasing $S$ token’s probability decreases editing faithfulness, while decreasing it and reallocating those values to the $\boldsymbol{T}$ and $E$ tokens increases editing faithfulness. Thus, for $\boldsymbol{A}_{\boldsymbol{O},j}$, we zero out $\boldsymbol{T}$ and $E$ tokens’ probabilities, reallocating them to $S$ to prevent edits outside the mask (blue dashed box in Fig. 4, bottom). In addtition, we redistribute the attention probability for the object region inside the mask by reducing the value of $S$ and redistributing it to $\boldsymbol{T}$ and $E$ (red dashed box in Fig. 4, bottom). We reduce $\boldsymbol{A}_{\boldsymbol{I},j=S}$ by $\lambda_{S}$, which decays linearly to 0 from $t=T$ to $t=1$, formulated as:
@@
where $W$ is a warm-up value that is linearly decayed from $\lambda$ to 0 in the early sampling steps $t<0.1T$ to increase the editing fidelity especially for small and difficult objects. Increasing $\lambda$ tends to enhance faithfulness but may introduce artifacts. We empirically found that $\lambda=0.5$ is the best trade-off between faithfulness and artifact-less results. Then, we redistribute $\lambda_{S}$ to $\boldsymbol{T}$ and $E$ with the ratios of $\lambda_{r}$ and $1-\lambda_{r}$, respectively. The attention probabilities, $\boldsymbol{A}_{\boldsymbol{I},j\in\boldsymbol{T}}$ and $\boldsymbol{A}_{\boldsymbol{I},j=E}$, of $\boldsymbol{T}$ and $E$ are updated as follows:
@@
where $N_{\boldsymbol{T}}$ denotes the number of text tokens. Increasing $\lambda_{r}$ may lead to enhancing the editing details for certain tokens, while decreasing it may increase the overall editing fidelity. $\lambda_{r}=0.5$ is empirically set for all experiments.
3.3 Disentangled Multi-instance Sampling
To disentangle the multi-instance VE process and reduce attention leakage, we propose our Disentangled Multi-instance Sampling (DMS). As shown in Fig. 3, DMS comprises of two sampling strategies: (1) Series Noise Sampling (SNS) shown in yellow box and (2) Parallel Noise Sampling (PNS) shown in blue box preceded by latent fusion (green box) and re-inversion (purple box). In the SNS, we independently edit each instance $i$ using its target caption $c_{i}$ and its mask $\boldsymbol{m_{i}}$ by first estimating each instance noise $\hat{n}_{i}^{1:N}$ as:
$\hat{n}_{i}^{1:N}=\epsilon_{\theta}(\{z_{t}^{1:N}\}_{i},c_{i},\boldsymbol{m_{i}},e,t),$ (4)
and the background noise $\tilde{n}^{1:N}$ using the inverted latents as:
$\tilde{n}^{1:N}=\epsilon_{\theta}(\tilde{z}_{t}^{1:N},c_{\varnothing},t).$ (5)
Then, we fuse the instance noise $\hat{n}_{i}^{1:N}$ with the inverted latent noise $\tilde{n}^{1:N}$ to obtain the edited instance noise $\{\bar{n}_{t-1}^{1:N}\}_{i}$ using masking as:
$\{\bar{n}_{t-1}^{1:N}\}_{i}=\hat{n}_{i}^{1:N}\cdot\boldsymbol{m_{i}}+\tilde{n}^{1:N}\cdot(1-\boldsymbol{m_{i}}).$ (6)
We then perform one DDIM [51] denoising step from $t\rightarrow t-1$ to obtain the instance latents $\{\hat{z}_{t-1}^{1:N}\}_{i}$ as:
$\{\hat{z}_{t-1}^{1:N}\}_{i}=\text{DDIM}(\{\hat{z}_{t}^{1:N}\}_{i},\{\bar{n}_{t-1}^{1:N}\}_{i},t).$ (7)
SNS requires the edited instance $i$ to appear within $\boldsymbol{m_{i}}$, as the background is replaced with the inverted latent noise at each step to ensure it remains unchanged. Our proposed IPR (Sec. 3.2) provides this necessary spatial control.
While we can perform multi-instance VE by solely using SNS until the end of denoising, the edited frames suffer from temporal inconsistency (see Tab. 3). Thus, we propose two approaches to mitigate this:
(i) we fuse the multiple instance latents $\{\dot{z}_{t}^{1:N}\}$ sampled independently from SNS as follows:
$\dot{z}_{t}^{1:N}=\Sigma_{i=1}^{M}(\{\hat{z}_{t}^{1:N}\}_{i}\cdot\boldsymbol{m_{i}})+\tilde{z}_{t}^{1:N}\cdot\boldsymbol{m_{B}},$ (8)
where $\boldsymbol{m_{B}}=\boldsymbol{1}-\Sigma_{i=1}^{M}\boldsymbol{m_{i}}$ denotes a background mask;
(ii) we propose re-inversion using DDIM for $l$ steps after the latent fusion to obtain $\ddot{z}_{t+l}^{1:N}$ followed by PNS from timestep $t+l\rightarrow t$ to yield the final fused latent $\ddot{z}_{t}^{1:N}$.
We then perform PNS for the remaining denoising steps. The goal of our PNS is to harmonize the independent instance latents obtained from the SNS while still decoupling the sampling using the individual target captions. In this sampling strategy, we use the re-inverted fused latents $\ddot{z}_{t}^{1:N}$ to estimate each instance noise $\hat{n}_{i}^{1:N}=\epsilon_{\theta}(\ddot{z}_{t}^{1:N},c_{i},\boldsymbol{m_{i}},e,t)$ using the instance caption $c_{i}$. We still estimate the noise $\tilde{n}^{1:N}=\epsilon_{\theta}(\tilde{z}_{t}^{1:N},c_{\varnothing},t)$ for the background using the inverted latents $\tilde{z}_{t}^{1:N}$ and an empty caption $c_{\varnothing}$. We then combine $\hat{n}_{i}^{1:N}$ and $\tilde{n}^{1:N}$ to a single noise through masking, and perform one DDIM step to obtain the latents $\hat{z}_{t-1}^{1:N}$ as:
$\hat{z}_{t-1}^{1:N}=\text{DDIM}(\hat{z}_{t}^{1:N},\Sigma_{i=1}^{M}(\hat{n}_{i}^{1:N}\cdot\boldsymbol{m_{i}})+\tilde{n}^{1:N}\cdot\boldsymbol{m_{B}},t).$ (9)
Finally, we decode $\hat{z}_{t-1}^{1:N}$ using the VAE decoder $\mathcal{D}$ to obtain the final edited frames $\hat{\boldsymbol{f}}=\mathcal{D}(\hat{z}_{0}^{1:N})$.